OpenStudy (anonymous):

Solve:

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 }x = 1\frac{ 5 }{ 7 }\]

OpenStudy (anonymous):

The number on right hand side is called as Mixed fraction... Mixed Fraction can be changed to normal fraction by using this: \[a \frac{b}{c} = \frac{a \cdot c + b}{c}\]

OpenStudy (anonymous):

So, firstly solve right hand side first.. Using above concept that I gave now: try to solve: \[1 \frac{5}{7 } = ??\]

OpenStudy (anonymous):

I am expecting a reply atleast...

OpenStudy (anonymous):

Sorry my computer was frozen @waterineyes

OpenStudy (anonymous):

12/7 @waterineyes

OpenStudy (anonymous):

@waterineyes

OpenStudy (anonymous):

You there??

OpenStudy (anonymous):

Sorry, this time my internet got frozen... :)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

It's okay and thank you for understanding

OpenStudy (anonymous):

So you have now: \[\frac{4}{3}x =\frac{12}{7}\]

OpenStudy (anonymous):

Hey, it is okay.. Can you cross multiply there and solve for x or I should tell now step by step???

OpenStudy (anonymous):

Or I think there is no need to cross multiply also here...

OpenStudy (anonymous):

um i think if i cross multiply ill come out with: 28/36 ? idk if i did it wrong

OpenStudy (anonymous):

First Step : Just multiply by 3 both the sides... And tell me what do you get.. Don't worry if it is wrong, I am here to correct you only.. :)

OpenStudy (anonymous):

okay, would i multiply 12/7 by 3? and get 36/21?

OpenStudy (anonymous):

You have to multiply by 3 both the sides, let us do one by one.. Multiply by 3 first on left hand side and tell me what did you get??

OpenStudy (anonymous):

And also, just multiply 3 with numerator part, don't multiply it with denominator part of left or right hand side.. Now do and show me where you have reached?

OpenStudy (anonymous):

ok i got \[\frac{ 12 }{ 3 }x = \frac{ 36 }{ 7 }\] ?

OpenStudy (anonymous):

Hey see, you cancel 3 with 3 on left hand side, you have written 12/3 that is 4 itself... See here: \[3 \times \frac{4}{3}x = 3 \times \frac{12}{7} \implies \cancel{3} \times \frac{4}{\cancel{3}}x = \frac{36}{7} \implies \color{green}{4x = \frac{36}{7}}\]

OpenStudy (anonymous):

For the purpose of cancellation, I have told you to multiply 3, BUT ONE THING YOU MUST REMEMBER, WHAT YOU ARE DOING ADDITIONALLY, YOU WILL DO WITH BOTH LEFT AS WELL AS RIGHT SIDE.. As, I am here multiplying additionally with 3, so I just can't multiply it only with left hand side, I have to multiply it with right hand side to maintain equality between LHS and RHS..

OpenStudy (anonymous):

Are you getting this??

OpenStudy (anonymous):

Now divide by 4 both the sides: \[\frac{4x}{4} = \frac{36}{4 \times 7} \implies x = \frac{\cancel{36}^9}{\cancel{4}^1 \times 7} \implies \color{blue}{x = \frac{9}{7}}\]

OpenStudy (anonymous):

I have to go now.. Please try to understand the last step what I have done, surely you will get.. If not, then when I will come online, do ask me where you are not getting... Going..

OpenStudy (anonymous):

ok thank you i get it! :) my computer is being really slow so sorry if this is really late

OpenStudy (anonymous):

@waterineyes ^